# Progressive Prompt Detailing for Improved Alignment in Text-to-Image Generative Models

### v1

@@ Abstract
Text-to-image generative models often struggle with long prompts detailing complex scenes, diverse objects with distinct visual characteristics and spatial relationships. In this work, we propose SCoPE (Scheduled interpolation of Coarse-to-fine Prompt Embeddings), a training-free method to improve text-to-image alignment by progressively refining the input prompt in a coarse-to-fine-grained manner. Given a detailed input prompt, we first decompose it into multiple sub-prompts which evolve from describing broad scene layout to highly intricate details. During inference, we interpolate between these sub-prompts and thus progressively introduce finer-grained details into the generated image. Our training-free plug-and-play approach significantly enhances prompt alignment, achieves an average improvement of up to +4% in Visual Question Answering (VQA) scores over the Stable Diffusion baselines on 85% of the prompts from the GenAI-Bench dataset.
@@ 1 Introduction
In this work, we propose SCoPE which stands for Scheduled interpolation of Coarse-to-fine Prompt Embeddings. SCoPE is a training-free approach that improves alignment between the provided (long) prompt and the generated image in diffusion models. Our key idea is to dynamically break down the input prompt into a series of sub-prompts starting from a coarse-grained description that captures the global scene layout to more fine-grained details. We draw inspiration from the findings in Park et al. [11] that diffusion denoising is a progressive coarse-to-fine generation process, where initial timesteps establish low-frequency, global structures, while later steps introduce high-frequency, fine-grained details. Specifically, while all prior methods rely on a single static embedding of the entire input prompt, SCoPE interpolates between progressively detailed prompt embeddings throughout the denoising process, thus generating the global scene layout before gradually introducing finer-grained details. We extensively evaluate SCoPE against several open-sourced models and show that SCoPE improves prompt alignment for long complex prompts (obtained from the GenAI-bench dataset [6]) (see Fig. 1) and achieves a +4% improvement in VQA-based text-image alignment scores over Stable Diffusion [17, 12] baselines. Notably, SCoPE is both training-free and easily extensible, requires minimal computational overhead of only $+0.7$ seconds per inference on $1$ A6000 GPU.
@@ 4 Experiments
SCoPE-v2-1 65.5 84.81% 34.9 77.56%
SD-v2-1 62.8 - 33.6 -
SCoPE-v1-4 65.1 82.00% 34.5 74.38%
SD-v1-4 62.8 - 33.3 -
SCoPE-XL 66.2 82.13% 35.3 65.56%
SDXL 63.8 - 34.6 -
@@
Quantitative results. As shown in Table 2, SCoPE consistently improves text-image alignment, achieving higher VQA Scores and CLIP scores compared to the baselines. For Stable Diffusion-2.1, SCoPE achieves a mean VQAScore of $\mathbf{65.5}$ across all 1600 prompts derived from GenAI-Bench, outperforming the baseline score of $62.8$. We also observe that SCoPE achieves an $\mathbf{84.81\%}$ win rate, indicating that in over $84\%$ of prompts, SCoPE-generated images were more aligned with the input text prompt (as measured by VQAScore). A similar trend is observed for CLIP Score, where SCoPE achieves a mean score of $\mathbf{34.9}$, surpassing the baseline score of $33.6$, with a $77.56\%$ win rate. For Stable Diffusion 1-4 and SDXL, SCoPE increases VQAScore to $\mathbf{65.1}$ and $\mathbf{66.2}$, with win rates of $82\%$ and $82.13\%$, respectively. In Table 4 we show that SCoPE outperforms the baseline model on varied prompt categories, such as “Spatial Relation,” “Counting,” “Negation,” etc.
@@ 6.2 Simplification Prompt
Basic Attribute 0.6612 0.6290 0.3521 0.3367 90.29% 82.22%
Scene 0.6600 0.6290 0.3490 0.3350 88.90% 79.80%
Spatial Relation 0.6594 0.6268 0.3535 0.3372 90.73% 82.19%
Action Relation 0.6626 0.6294 0.3549 0.3387 90.82% 82.46%
Part Relation 0.6646 0.6316 0.3572 0.3412 88.86% 81.66%
Advanced Counting 0.6578 0.6235 0.3501 0.3340 90.27% 83.19%
Comparison 0.6579 0.6268 0.3505 0.3351 91.36% 80.25%
Differentiation 0.6615 0.6271 0.3489 0.3342 91.22% 82.43%
Negation 0.6632 0.6302 0.3483 0.3326 91.35% 80.69%
Universal 0.6526 0.6185 0.3451 0.3302 92.52% 79.59%
Table 3: VQA and CLIP Scores on different tags from the GenAI-Bench [6] dataset for SCoPE and Stable Diffusion-2.1 (SD-v2-1).

### v2

@@ Abstract
Text-to-image generative models often struggle with long prompts detailing complex scenes, diverse objects with distinct visual characteristics and spatial relationships. In this work, we propose SCoPE (Scheduled interpolation of Coarse-to-fine Prompt Embeddings), a training-free method to improve text-to-image alignment by progressively refining the input prompt in a coarse-to-fine-grained manner. Given a detailed input prompt, we first decompose it into multiple sub-prompts which evolve from describing broad scene layout to highly intricate details. During inference, we interpolate between these sub-prompts and thus progressively introduce finer-grained details into the generated image. Our training-free plug-and-play approach significantly enhances prompt alignment, achieves an average improvement of more than +8 in Visual Question Answering (VQA) scores over the Stable Diffusion baselines on 83% of the prompts from the GenAI-Bench dataset.
@@ 1 Introduction
In this work, we propose SCoPE which stands for Scheduled interpolation of Coarse-to-fine Prompt Embeddings. SCoPE is a training-free approach that improves alignment between the provided (long) prompt and the generated image in diffusion models. Our key idea is to dynamically break down the input prompt into a series of sub-prompts starting from a coarse-grained description that captures the global scene layout to more fine-grained details. We draw inspiration from the findings in Park et al. [11] that diffusion denoising is a progressive coarse-to-fine generation process, where initial timesteps establish low-frequency, global structures, while later steps introduce high-frequency, fine-grained details. Specifically, while all prior methods rely on a single static embedding of the entire input prompt, SCoPE interpolates between progressively detailed prompt embeddings throughout the denoising process, thus generating the global scene layout before gradually introducing finer-grained details. We extensively evaluate SCoPE against several open-sourced models and show that SCoPE improves prompt alignment for long complex prompts (obtained from the GenAI-bench dataset [6]) (see Fig. 1) and achieves a +8 improvement in VQA-based text-image alignment scores over Stable Diffusion [17, 12] baselines. Notably, SCoPE is both training-free and easily extensible, requires minimal computational overhead of only $+0.7$ seconds per inference on $1$ A6000 GPU.
@@ 4 Experiments
SCoPE-v2-1 87.3 83.88% 34.9 77.56%
SD-v2-1 79.2 - 33.6 -
SCoPE-v1-4 84.7 83.44% 34.5 74.38%
SD-v1-4 76.6 - 33.3 -
SCoPE-XL 87.7 73.00% 35.3 65.56%
SDXL 82.9 - 34.6 -
@@
Quantitative results. As shown in Table 2, SCoPE consistently improves text-image alignment, achieving higher VQA Scores and CLIP scores compared to the baselines. For Stable Diffusion-2.1, SCoPE achieves a mean VQAScore of $\mathbf{87.3}$ across all 1600 prompts derived from GenAI-Bench, outperforming the baseline score of $79.2$. We also observe that SCoPE achieves an $\mathbf{83.88\%}$ win rate, indicating that in over $83\%$ of prompts, SCoPE-generated images were more aligned with the input text prompt (as measured by VQAScore). A similar trend is observed for CLIP Score, where SCoPE achieves a mean score of $\mathbf{34.9}$, surpassing the baseline score of $33.6$, with a $77.56\%$ win rate. For Stable Diffusion 1-4 and SDXL, SCoPE increases VQAScore to $\mathbf{84.7}$ and $\mathbf{87.7}$, with win rates of $83.44\%$ and $73.00\%$, respectively. In Table 4 we show that SCoPE outperforms the baseline model on varied prompt categories, such as “Spatial Relation,” “Counting,” “Negation,” etc.
@@ 6.2 Simplification Prompt
Basic Attribute 0.8683 0.7878 0.3521 0.3367 84.53% 82.22%
Scene 0.8857 0.8145 0.3490 0.3350 83.78% 79.80%
Spatial Relation 0.8661 0.7783 0.3535 0.3372 83.87% 82.19%
Action Relation 0.8649 0.7675 0.3549 0.3387 83.51% 82.46%
Part Relation 0.8678 0.7820 0.3572 0.3412 86.46% 81.66%
Advanced Counting 0.8743 0.7949 0.3501 0.3340 84.66% 83.19%
Comparison 0.8632 0.7845 0.3505 0.3351 82.10% 80.25%
Differentiation 0.8472 0.7687 0.3489 0.3342 82.43% 82.43%
Negation 0.8780 0.8105 0.3483 0.3326 86.17% 80.69%
Universal 0.8997 0.8478 0.3451 0.3302 82.31% 79.59%
Table 3: VQA and CLIP Scores on different tags from the GenAI-Bench [6] dataset for SCoPE and Stable Diffusion-2.1 (SD-v2-1). Absolute improvements are consistent across all categories.